# Dark Energy Survey Year 6 Results: Cell-based Coadds and Metadetection Weak Lensing Shape Catalogue

### v1

@@ 4.6 Summary Statistics of the Catalogue
Here we present summary statistics for the catalogue after the selections described in Sec. 4.4. As we have described earlier, for Metadetection we compute the shear response averaged over the catalogue rather than each object using Eqn. 6. The shear response over the catalogue is $\langle R\rangle=0.817$. The magnitude limit of our catalogue is $\sim$ 0.5 magnitude shallower than that of Gold catalogue (Bechtol et al., 2025) in $i$-band, resulting in $m_{i}\sim 22.8$. This number very approximately represents the median depth of the catalogue across the survey at a S/N10similar-toabsent10$\sim$ 10∼ 10.
We measure the statistical power of the catalogue by computing the standard error of the mean shear across the catalog, $\sigma_{\rm\gamma}=\sigma_{\rm e}/\sqrt{N_{\rm eff}}$. Following the definitions of Heymans et al. 2012 (H12) and Chang et al. 2013 (C13), the shape noise ($\sigma_{\rm e}$; standard deviation of intrinsic ellipticities) and the effective number density ($n_{\rm eff}$) are computed as follows. For the C13 definition, we compute
@@
We report the effective galaxy number density and the variance of the estimated shear of the objects in our Metadetection shape catalogue in Table 3. One finds that the shape noise in Y6 has increased compared to Y3. We speculate that this increase is due to detecting fainter objects than Y3 and possibly detecting more early-type galaxies that have larger intrinsic scatter. The effective number density increase between Y3 and Y6 is difficult to predict because it depends on the faint-end slope of the luminosity function. However, note that we lose approximately two observations of each object during the cell-coadding process, so that the number density increased less than it might have otherwise. Figure 6 shows the weighted number density and shape noise in the H12 definition. The statistical power of the Y6 shape catalogue has increased by 12% for H12 definition and 10% for C13 definition, compared to that of Y3 weak lensing shape catalogue (Gatti, Sheldon et al., 2021).
@@
Y6 7.97, 8.22 0.270, 0.289 $+1.94\times 10^{-4}$ $-0.314\times 10^{-4}$
Y3 5.32, 5.59 0.255, 0.261 $+3.52\times 10^{-4}$ $+0.594\times 10^{-4}$
Figure 7 shows the Y6 catalogue precision compared to other DES weak lensing shape catalogue releases – Science Verification (SV, Jarvis et al. 2016), Year 1 (Y1, Zuntz et al. 2018), Year 3 (Y3, Gatti, Sheldon et al. 2021). The figure presents three key quantities, the survey area, the shear information density, and total shear information content. The histogram shows the number density weighted by the shape noise2 in HEALPix pixels with nside=1024, which we call shear information density per unit area. The total shear information is approximately the survey area $\times$ shear information density. The inverse of the square root of this number is the error on the mean shear across the survey. The figure shows several major achievements by the DES collaboration. The survey has increased in area substantially, increased in depth (increasing the shear information density), and become more homogeneous. These changes are a testament to the hard work of dozens of scientists working on weak lensing shape measurement in the DES collaboration over more than a decade.
Figure 7: A histogram showing the shear information density ($n_{\rm eff}/\sigma^{2}_{e}$) in HEALPix pixels (nside=1024) for the DES Y6 Metadetection catalogue footprint, compared with the same quantity from several catalogue releases from the Dark Energy Survey – Science Verification (SV, Jarvis et al. 2016), Year 1 (Y1, Zuntz, Sheldon et al. 2018), Year 3 (Y3, Gatti, Sheldon et al. 2021). Each histogram has the same area under the curve, and the median of each histogram is shown as the dotted line for each release. The interquartile range (IQR) for each release is $\{38.22,38.57,37.55,36.81\}$ for SV through Y6. The histogram is accompanied by the panels showing the survey area (left) and the total shear information content of the catalogue (right), which is the sum of pixel areas for each survey times the shear information density. We can see that as the DES survey has progressed, the data set has become more uniform and its total information content has increased substantially.
@@ 5.1 Mean Shear as a Function of PSF Properties
Figure 8: Mean shear ($e_{1}$, $e_{2}$) as a function of several PSF and galaxy properties. The mean shear is corrected by the shear response and the global mean shear in the catalogue is subtracted from each data point. The linear fit (slope and offset) to the data points was computed with ordinary least-squares. The error bars were estimated by dividing the sample into 200 patches in the sky and computed via jackknife covariance estimation. Top-Left: as a function of PSF shape. Middle-Left: as a function of PSF size. Bottom-Left: as a function of galaxy size. Top-Right: as a function of galaxy and PSF size ratio. Middle-Right: as a function of galaxy $g-i$ color. Bottom-Right: as a function of galaxy signal-to-noise (S/N).
Figure 9: Mean shear (top row: $e_{1}$ and bottom row: $e_{2}$) as a function of PSF shapes where the catalogue is split in 3 ($g-i$) color bins. Left column: color range is set between $-2.00<g-i<0.76$. Middle column: color range is set between $1.490.76<g-i<$ 1.49. Right column: color range is set between $4.001.49<g-i<$ 4.00. Ordinary least-square fits are performed on the data points to produce the black line, and red dotted line for the linear fit performed on the Y3 shape catalogue.
Table 4: Comparison of mean shear ($e_{1}$ and $e_{2}$) of the objects in Y3 and Y6 shape catalogue. We split the objects into three color ($g-i$) ranges (blue: [-2.00, 0.76], mid: [0.76, 1.49], red: [1.49, 4.00]). For the Y6 data, we compute the shear response of the objects that are selected to be in each defined color range. For Y3, we used shapes from the Y3 catalogue, and the shear and selection responses were re-computed for each color sample. We find improved additive bias control as a function of color.
@@
$\alpha^{(2)}$ $0.0104_{-0.0100}^{+0.0100}$ $0.00682_{-0.00796}^{+0.00801}$
$\beta^{(2)}$ $0.737_{-0.570}^{+0.566}$ $1.19_{-0.23}^{+0.23}$
$\eta^{(2)}$ $2.44_{-7.11}^{+7.04}$ $-1.46_{-1.37}^{+1.38}$
$\alpha^{(4)}$ $-0.0128_{-0.0292}^{+0.0290}$ $-0.00747_{-0.02491}^{+0.02488}$
$\beta^{(4)}$ $1.95_{-0.272}^{+0.274}$ $2.14_{-0.21}^{+0.21}$
$\eta^{(4)}$ $3.15_{-2.59}^{+2.60}$ $0.461_{-0.703}^{+0.702}$
$\eta^{(24)}$ $-1.28_{-0.664}^{+0.663}$ $-0.103_{-0.100}^{+0.100}$
$\eta^{(42)}$ $-10.8_{-18.9}^{+18.9}$ $-0.646_{-1.728}^{+1.754}$
@@ 5.2 Additive Biases Due to PSF Leakage and Modeling errors
Following Paulin-Henriksson et al. 2008; Rowe 2010; Jarvis et al. 2016, we can decompose $\delta e^{\rm sys}_{\rm PSF}$ into terms that scale like the PSF model shape, the PSF star image shape, and the product of the PSF size residual times the PSF star shape. Recently, Zhang et al. (2023b, a) pointed out that higher-order radial moments that scale as the radius to the fourth power, as opposed to the radius squared, are also important in this expansion and such terms were used the HSC Y3 analysis (Li et al., 2022). We follow Zhang et al. (2023a) and include spin-0 and spin-2 fourth-order moments as well,
$e^{\rm(4)}_{\rm PSF}=\frac{M_{31}+iM_{13}}{T^{2}}-3e^{\rm(2)}_{\rm PSF},T^{\rm(4)}_{\rm PSF}=\frac{M_{22}}{M_{11}}.$ (16)
@@
$\delta e^{\rm sys}_{\rm PSF}=\alpha^{(2)}e_{\rm PSF}+\beta^{(2)}\Delta e_{\rm PSF}+\eta^{(2)}(e_{*}\Delta T_{\rm PSF}/T_{*})$
$+\alpha^{(4)}e^{(4)}_{\rm PSF}+\beta^{(4)}\Delta e^{(4)}_{\rm PSF}+\eta^{(4)}(e^{(4)}_{*}\Delta T^{(4)}_{\rm PSF}/T^{(4)}_{*})$
$+\eta^{(24)}(e_{*}\Delta T^{(4)}_{\rm PSF}/T^{(4)}_{*})+\eta^{(42)}(e^{(4)}_{*}\Delta T_{\rm PSF}/T_{*})$ (17)
@@
To constrain this model, we measure the full set of PSF shape auto-correlations and the full set of PSF-galaxy shape cross-correlations. Using the model above, we can then make predictions for these signals and constrain the model coefficients. See Appendix D, Eqn. 22-29, for the full expressions. The galaxy-shear cross-correlation statistics on the left-hand side of Eqn. 22-29 are usually denoted as $\tau$-statistics. The auto-correlations of the PSF quantities are usually denoted as $\rho$-statistics. Schutt et al. (2025) present the $\rho$-statistics for the Y6 PSF models and demonstrate improvements relative to the DES Y3 PSF models (see their Fig. 20). In our case, we evaluate the PSF models at the locations of the reserve stars using the median galaxy color, producing the reserve star catalogue. We then use this catalog, combined with shape measurements performed on the stars directly and the Metadetection catalog, to compute the various two-point correlation functions in 32 angular bins ranging between 2.5 - 250 arcmin. We estimate the covariance of our measurements from the 800 mock catalogues (see Sec. 2.2), applying statistical corrections for the finite number of mock realizations (Percival et al., 2022).
@@ 6.2 Tangential Shear Around Cell-based Coadd, Coadd Tile, and Field/Exposure Centers
In order to better understand the source of this signal, we split our galaxy sample by their $g-i$ colors. Figure 17 shows the same measurement of tangential shear in different coordinate systems. We again find no trend around cell-based coadd and coadd centers, but find a significant deviation from the null signal for the field/exposure centers around 7-8 arcmin, 10-20 arcmin, and 40-60 arcmin scale in the right panel of Fig 17. This deviation corresponds to the deviations seen in the right panel of Fig 16. Some of these effects can be understood as PSF modeling errors due to Metadetection using the median galaxy color for the PSF. Figure 18 shows the same statistics as we measured for the galaxies, but for the PSF modeling residuals ($e_{*}-e_{\rm PSF}$) when using PSF models evaluated at the median galaxy color versus the stars actual color. Even when the PSFs are evaluated at their actual star color, the right panel of the figure suggests the non-zero patterns at scales around 10-20 arcmin, which corresponds to the scales of each CCD.
@@
Figure 17: Tangential shear around cell-based coadd, coadd, and field/exposure centers where the galaxy sample is split by their measured $g-i$ color (blue: [-2.00, 0.76], mid: [0.76, 1.49], red: [1.49, 4.00]).
Figure 18: Tangential projection of residual PSF shapes after model subtraction around field/exposure centers where the star sample is split by their measured $g-i$ color (blue: [-2.00, 0.76], mid: [0.76, 1.49], red: [1.49, 4.00]). Left panel shows the case when the PSFs were drawn at their median color, which is the case in our Metadetection shape catalogue. Right panel shows the case when the PSFs were drawn at their actual star colors.
@@ 6.4 Null Tests with Survey Property Maps
  • Dust Extinction – the dust extinction map in Schlegel et al. (1998) from which we used $E(B-V)$ values to correct for the reddening. (nside=4096)
@@
Figure 20: Statistical significance of the dependence of the shear measurements on survey properties. The horizontal axis gives the statistical significance of the slope of the best-fit linear relation between the shear signals and the survey property. $E1$/$E2$ in the y-axis denotes the $e_{1}$/$e_{2}$ component of the shear correlated with the survey property. The significance of the observed slope is computed using 800 Cosmogrid simulations.
@@ 7 Image Simulation-based Tests
So far, we have discussed the catalogue-level systematics tests and mostly examined the sources of additive bias, which we can calibrate from data. The calibration of multiplicative bias, however, needs to be characterized and carried out in realistic image simulations. For this purpose, we use an image simulation suite generated using an update version of the DES Y3 image simulation code and data (MacCrann, Becker et al., 2022). In this work, we present a short description of the simulations and defer a more detailed presentation to Mau, Becker et al. 2025.
The image simulations are generated independently for each coadd tile. We use a set of $\sim 1000$ randomly selected tiles. For each coadd tile, we simulate each input single-epoch (CCD) image using the observed pixel masks, backgrounds, WCS, and PSF solutions. The input single-epoch images are then put through the identical cell-based coadding and Metadetection pipelines as the actual DES data. The input catalogue of galaxies for the simulations is derived from the same catalogue used in MacCrann, Becker et al. (2022), but with updates to include realistic galaxy clustering from the Cardinal simulation (To et al., 2024) and corrections to the bright end of the magnitude function for the deficit of bright objects in the original catalogue. We again used fitvd [^24] https://github.com/esheldon/fitvd models fit to the data as the input sources (MacCrann, Becker et al., 2022). For the stars, we used a combination of the LSST trilegal simulations (Dal Tio et al., 2022) and Gaia EDR3 (Gaia Collaboration et al., 2021) catalogues. This combination allows us to place very bright stars in their correct locations. Our Metadetection pipeline uses the locations of Gaia stars for special processing steps and so this detail is important. When simulating objects, we use color and position of the object to evaluate the PSF model, allowing our simulations to test both the effects of using the median galaxy color in the Metadetection code and the effects of ignoring the PSF spatial dependence within single cell-based coadds.
@@
    The difference between median_color and fiducial configurations validates that our ability to recover shear in settings with realistic levels of blending is robust even with the approach of evaluating PSF models at median galaxy color. Using the paired jacknife technique, we find $\Delta m=(0.035\pm 2.017)\times 10^{-3}$, $\Delta c_{1}=(-3.5\pm 2.5)\times 10^{-5}$ and $\Delta c_{2}=(-7.2\pm 2.5)\times 10^{-5}$.
@@
A final calibration of $m$ and prior/uncertainty for the cosmological analyses, including the effects of blending due to projected sources between different redshifts, will be derived and reported in Mau, Becker et al. 2025.
Table 7: Measurements of multiplicative and additive shear biases via jackknifing Eqn. 18 over measurements from simulated images of Y6 tiles. The reported uncertainties correspond to 3 standard errors of the mean.
Configuration $m_{1}$ [$10^{-3}$] $c_{2}$ [$10^{-4}$] ##\## tiles
grid-exp $-0.27\pm 0.10$ $+0.002\pm 0.018$ 97
grid $-3.8\pm 10$ $-0.5\pm 3.7$ 395
grid-median_color $-4.3\pm 9.9$ $+0.1\pm 3.7$ 395
no_stars $+3.5\pm 9.5$ $+0.4\pm 4.1$ 395
median_color $+3.4\pm 9.8$ $+0.8\pm 4.1$ 395
fiducial $+3.4\pm 6.1$ $+1.0\pm 2.6$ 981

### v2

@@ 4.6 Summary Statistics of the Catalogue
Here we present summary statistics for the catalogue after the selections described in Sec. 4.4. As we have described earlier, for Metadetection we compute the shear response averaged over the catalogue rather than each object using Eqn. 6. The shear response over the catalogue is $\langle R\rangle=0.817$. We measure the statistical power of the catalogue by computing the standard error of the mean shear across the catalog, $\sigma_{\rm\gamma}=\sigma_{\rm e}/\sqrt{N_{\rm eff}}$. Following the definitions of Heymans et al. 2012 (H12) and Chang et al. 2013 (C13), the shape noise ($\sigma_{\rm e}$; standard deviation of intrinsic ellipticities) and the effective number density ($n_{\rm eff}$) are computed as follows. For the C13 definition, we compute
@@
We report the effective galaxy number density and the variance of the estimated shear of the objects in our Metadetection shape catalogue in Table 3. One finds that the shape noise in Y6 has increased compared to Y3. We speculate that this increase is due to detecting fainter objects than Y3 and possibly detecting more early-type galaxies that have larger intrinsic scatter. The effective number density increase between Y3 and Y6 is difficult to predict because it depends on the faint-end slope of the luminosity function. However, note that we lose approximately two observations of each object during the cell-coadding process, so that the number density increased less than it might have otherwise. Figure 6 shows the weighted number density and shape noise in the H12 definition. The statistical power of the Y6 shape catalogue has increased by 12% for H12 definition and 10% for C13 definition, compared to that of Y3 weak lensing shape catalogue (Gatti, Sheldon, and others, 2021).
@@
Y6 7.97, 8.22 0.270, 0.289 $1.94\times 10^{-4}$ $0.314\times 10^{-4}$
Y3 5.32, 5.59 0.255, 0.261 $3.52\times 10^{-4} +$$0.594\times 10^{-4}$
Figure 7 shows the Y6 catalogue precision compared to other DES weak lensing shape catalogue releases – Science Verification (SV, Jarvis et al. 2016), Year 1 (Y1, Zuntz et al. 2018), Year 3 (Y3, Gatti, Sheldon, and others 2021). The figure presents three key quantities, the survey area, the shear information density, and total shear information content. The histogram shows the number density weighted by the shape noise2 in HEALPix pixels with nside=1024, which we call shear information density per unit area. The total shear information is approximately the survey area $\times$ shear information density. The inverse of the square root of this number is the error on the mean shear across the survey. The figure shows several major achievements by the DES collaboration. The survey has increased in area substantially, increased in depth (increasing the shear information density), and become more homogeneous. These changes are a testament to the hard work of dozens of scientists working on weak lensing shape measurement in the DES collaboration over more than a decade.
Figure 7: A histogram showing the shear information density ($n_{\rm eff}/\sigma^{2}_{e}$) in HEALPix pixels (nside=1024) for the DES Y6 Metadetection catalogue footprint, compared with the same quantity from several catalogue releases from the Dark Energy Survey – Science Verification (SV, Jarvis et al. 2016), Year 1 (Y1, Zuntz, Sheldon, and others 2018), Year 3 (Y3, Gatti, Sheldon, and others 2021). Each histogram has the same area under the curve, and the median of each histogram is shown as the dotted line for each release. The interquartile range (IQR) for each release is $\{38.22,38.57,37.55,36.81\}$ for SV through Y6. The histogram is accompanied by the panels showing the survey area (left) and the total shear information content of the catalogue (right), which is the sum of pixel areas for each survey times the shear information density. We can see that as the DES survey has progressed, the data set has become more uniform and its total information content has increased substantially.
@@ 5.1 Mean Shear as a Function of PSF Properties
Figure 8: Mean shear ($e_{1}$, $e_{2}$) as a function of several PSF and galaxy properties. The mean shear is corrected by the shear response and the global mean shear in the catalogue is subtracted from each data point. The linear fit (slope and offset) to the data points was computed with ordinary least-squares. The error bars were estimated by dividing the sample into 200 patches in the sky and computed via jackknife covariance estimation. Top-Left: as a function of PSF shape. Middle-Left: as a function of PSF size. Bottom-Left: as a function of galaxy size. Top-Right: as a function of galaxy and PSF size ratio. Middle-Right: as a function of galaxy gig-i color. Bottom-Right: as a function of galaxy signal-to-noise (S/N).
Figure 9: Mean shear (top row: $e_{1}$ and bottom row: $e_{2}$) as a function of PSF shapes where the catalogue is split in 3 (gig-i) color bins. Left column: color range is set between 2.00<gi<0.76-2.00<g-i<0.76. Middle column: color range is set between 0.76<gi<1.490.76<g-i<1.49. Right column: color range is set between 1.49<gi<4.001.49<g-i<4.00. Ordinary least-square fits are performed on the data points to produce the black line, and red dotted line for the linear fit performed on the Y3 shape catalogue.
Table 4: Comparison of mean shear ($e_{1}$ and $e_{2}$) of the objects in Y3 and Y6 shape catalogue. We split the objects into three color (gig-i) ranges (blue: [-2.00, 0.76], mid: [0.76, 1.49], red: [1.49, 4.00]). For the Y6 data, we compute the shear response of the objects that are selected to be in each defined color range. For Y3, we used shapes from the Y3 catalogue, and the shear and selection responses were re-computed for each color sample. We find improved additive bias control as a function of color.
@@
$\alpha^{(2)}$ $00980_{-0.00954}^{+0.0108} 0.0$$00682_{-0.00796}^{+0.00801}$
$\beta^{(2)}$ $0.737_{-0.571}^{+0.576}$ $1.19_{-0.23}^{+0.23}$
$\eta^{(2)}$ $2.71_{-7.47}^{+6.90}$ $1.46_{-1.37}^{+1.38}$
$\alpha^{(4)}$ $0115_{-0.0307}^{+0.0283} 0.0$$00747_{-0.02491}^{+0.02488}$
$\beta^{(4)}$ $1.94_{-0.268}^{+0.285}$ $2.14_{-0.21}^{+0.21}$
$\eta^{(4)}$ $3.10_{-2.60}^{+2.67}$ $0.461_{-0.703}^{+0.702}$
$\eta^{(24)}$ $1.27_{-0.677}^{+0.675}$ $0.103_{-0.100}^{+0.100}$
$\eta^{(42)}$ $11.8_{-18.3}^{+20.2}$ $0.646_{-1.728}^{+1.754}$
@@ 5.2 Additive Biases Due to PSF Leakage and Modeling errors
Following Paulin-Henriksson et al. 2008; Rowe 2010; Jarvis et al. 2016, we can decompose $\delta e^{\rm sys}_{\rm PSF}$ into terms that scale like the PSF model shape, the PSF star image shape, and the product of the PSF size residual times the PSF star shape. Recently, Zhang et al. (2023b, a) pointed out that higher-order radial moments that scale as the radius to the fourth power, as opposed to the radius squared, are also important in this expansion and such terms were used the HSC Y3 analysis (Li et al., 2022). We follow Zhang et al. (2023a) and include spin-0 and spin-2 fourth-order moments as well (see Sec. 4 of Schutt et al. (2025) for the definitions of the PSF moments).
@@
$\delta e^{\rm sys}_{\rm PSF}=\alpha^{(2)}e_{\rm PSF}+\beta^{(2)}\Delta e_{\rm PSF}+\eta^{(2)}(e_{*}\Delta T_{\rm PSF}/T_{\rm PSF})$
$+\alpha^{(4)}e^{(4)}_{\rm PSF}+\beta^{(4)}\Delta e^{(4)}_{\rm PSF}+\eta^{(4)}(e^{(4)}_{*}\Delta T^{(4)}_{\rm PSF}/T^{(4)}_{\rm PSF})$
$+\eta^{(24)}(e_{*}\Delta T^{(4)}_{\rm PSF}/T^{(4)}_{\rm PSF})+\eta^{(42)}(e^{(4)}_{*}\Delta T_{\rm PSF}/T_{\rm PSF})$ (16)
@@
To constrain this model, we measure the full set of PSF shape auto-correlations and the full set of PSF-galaxy shape cross-correlations. Using the model above, we can then make predictions for these signals and constrain the model coefficients. See Appendix D, Eqn. 21-28, for the full expressions. The galaxy-shear cross-correlation statistics on the left-hand side of Eqn. 21-28 are usually denoted as $\tau$-statistics. The auto-correlations of the PSF quantities are usually denoted as $\rho$-statistics. Schutt et al. (2025) present the $\rho$-statistics for the Y6 PSF models and demonstrate improvements relative to the DES Y3 PSF models (see their Fig. 20). In our case, we evaluate the PSF models at the locations of the reserve stars using the median galaxy color, producing the reserve star catalogue. We then use this catalog, combined with shape measurements performed on the stars directly and the Metadetection catalog, to compute the various two-point correlation functions in 32 angular bins ranging between 2.5 - 250 arcmin. We estimate the covariance of our measurements from the 800 mock catalogues (see Sec. 2.2), applying statistical corrections for the finite number of mock realizations (Percival et al., 2022).
@@ 6.2 Tangential Shear Around Cell-based Coadd, Coadd Tile, and Field/Exposure Centers
In order to better understand the source of this signal, we split our galaxy sample by their gig-i colors. Figure 17 shows the same measurement of tangential shear in different coordinate systems. We again find no trend around cell-based coadd and coadd centers, but find a significant deviation from the null signal for the field/exposure centers around 7-8 arcmin, 10-20 arcmin, and 40-60 arcmin scale in the right panel of Fig 17. This deviation corresponds to the deviations seen in the right panel of Fig 16. Some of these effects can be understood as PSF modeling errors due to Metadetection using the median galaxy color for the PSF. Figure 18 shows the same statistics as we measured for the galaxies, but for the PSF modeling residuals ($e_{*}-e_{\rm PSF}$) when using PSF models evaluated at the median galaxy color versus the stars actual color. Even when the PSFs are evaluated at their actual star color, the right panel of the figure suggests the non-zero patterns at scales around 10-20 arcmin, which corresponds to the scales of each CCD.
@@
Figure 17: Tangential shear around cell-based coadd, coadd, and field/exposure centers where the galaxy sample is split by their measured gig-i color (blue: [-2.00, 0.76], mid: [0.76, 1.49], red: [1.49, 4.00]).
Figure 18: Tangential projection of residual PSF 2nd order (top) and 4th order (bottom) shapes around field/exposure centers where the star sample is split by their measured gig-i color (blue: [-2.00, 0.76], mid: [0.76, 1.49], red: [1.49, 4.00]). Left panel shows the case when the PSFs were drawn at their median color, which is the case in our Metadetection shape catalogue. Right panel shows the case when the PSFs were drawn at their actual star colors.
@@ 6.4 Null Tests with Survey Property Maps
  • Dust Extinction – the dust extinction map in Schlegel et al. (1998) from which we used E(BV)E(B-V) values to correct for the reddening. (nside=4096)
@@
Figure 20: Statistical significance of the dependence of the shear measurements on survey properties. The horizontal axis gives the statistical significance of the slope of the best-fit linear relation between the shear signals and the survey property. E1E1/$E2$ in the y-axis denotes the $e_{1}$$e_{2}$ component of the shear correlated with the survey property. The significance of the observed slope is computed using 800 Cosmogrid simulations.
@@ 7 Image Simulation-based Tests
So far, we have discussed the catalogue-level systematics tests and mostly examined the sources of additive bias, which we can calibrate from data. The calibration of multiplicative bias, however, needs to be characterized and carried out in realistic image simulations. For this purpose, we use an image simulation suite generated using an update version of the DES Y3 image simulation code and data (MacCrann, Becker et al., 2022). In this work, we present a short description of the simulations and defer a more detailed presentation to Mau, Becker, and others 2025.
The image simulations are generated independently for each coadd tile. We use a set of $\sim 1000$ randomly selected tiles. For each coadd tile, we simulate each input single-epoch (CCD) image using the observed pixel masks, backgrounds, WCS, and PSF solutions. [^24] For the image simulations, we make the simulation and run the analysis using a simpler WCS from SCAMP (Bertin et al., 2002) for each CCD image. We made this change so that we could fully integrate the analysis pipeline with that used for synthetic-source injection (Anbajagane et al., 2025). The input single-epoch images are then put through the identical cell-based coadding and Metadetection pipelines as the actual DES data. The input catalogue of galaxies for the simulations is derived from the same catalogue used in MacCrann, Becker et al. (2022), but with updates to include realistic galaxy clustering from the Cardinal simulation (To et al., 2024) and corrections to the bright end of the magnitude function for the deficit of bright objects in the original catalogue. We again used fitvd [^25] https://github.com/esheldon/fitvd models fit to the data as the input sources (MacCrann, Becker et al., 2022). For the stars, we used a combination of the LSST trilegal simulations (Dal Tio et al., 2022) and Gaia EDR3 (Gaia Collaboration et al., 2021) catalogues. This combination allows us to place very bright stars in their correct locations. Our Metadetection pipeline uses the locations of Gaia stars for special processing steps and so this detail is important. When simulating objects, we use color and position of the object to evaluate the PSF model, allowing our simulations to test both the effects of using the median galaxy color in the Metadetection code and the effects of ignoring the PSF spatial dependence within single cell-based coadds.
@@
    The difference between median_color and fiducial configurations validates that our ability to recover shear in settings with realistic levels of blending is robust even with the approach of evaluating PSF models at median galaxy color. Using the paired jackknife technique, we find $\Delta m=(0.035\pm 2.017)\times 10^{-3}$, $\Delta c_{1}=(-3.5\pm 2.5)\times 10^{-5}$ and $\Delta c_{2}=(-7.2\pm 2.5)\times 10^{-5}$.
@@
A final calibration of $m$ and prior/uncertainty for the cosmological analyses, including the effects of blending due to projected sources between different redshifts, will be derived and reported in Mau, Becker, and others 2025.
Table 7: Measurements of multiplicative and additive shear biases via jackknifing Eqn. 17 over measurements from simulated images of Y6 tiles. The reported uncertainties correspond to 3 standard errors of the mean.
Configuration $m_{1}$ $10^{-3}$] $c_{2}$ $10^{-4}$] #\# tiles
grid-exp $27\pm 0.10 +0.0$$002\pm 0.018$ 97
grid $3.8\pm 10$ $0.5\pm 3.7$ 395
grid-median_color $4.3\pm 9.9$ $0.1\pm 3.7$ 395
no_stars $3.5\pm 9.5 +$$0.4\pm 4.1$ 395
median_color $3.4\pm 9.8 +$$0.8\pm 4.1$ 395
fiducial $3.4\pm 6.1 +$$1.0\pm 2.6$ 981